OpenStudy (anonymous):

Solve 3^(2x) = 7^(x−1).

OpenStudy (anonymous):

hellpp

OpenStudy (jdoe0001):

\(\bf 3^{2x} = 7^{x-1}\) right? for I don't see it happening

OpenStudy (anonymous):

yes @jdoe0001

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

hmm hold the mayo

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (jdoe0001):

hmmm I assume you've covered logarithms?

OpenStudy (anonymous):

First step: Take the logarithm of both sides.

OpenStudy (anonymous):

ao log 3^(2x)= log 7^(x-1)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

OK. When you have a power, the way to take the log is as follows:\[\log a ^{b} = b \log a\]For example\[\log 5^\left( x+2 \right) = \left( x+2 \right) \log 5\] Try this with your question

OpenStudy (anonymous):

3^(2x) log= 7^(x-1)= log

OpenStudy (anonymous):

\[3^{2x}=7^{x-1}=7^x*7^{-1}\] \[\left( 3^2 \right)^x=\frac{ 7^x }{ 7 }\] \[\frac{ 9^x }{ 7^x }=\frac{ 1 }{ 7 }\] \[\left( \frac{ 9 }{ 7 } \right)^x=\frac{ 1 }{ 7 }\] take log and find x

OpenStudy (anonymous):

−7.74293 7.74293 −1 1 These are the answer choices <3

OpenStudy (anonymous):

Not quite. Let's look at just the left hand side. You are trying to determine\[\log 3^{2x} = ?\]Compare that with the general case\[\log a^b = b \log a\]Substitute a = 3 and b = 2x. What do you get?

OpenStudy (anonymous):

2x log 3

OpenStudy (anonymous):

was i right

OpenStudy (anonymous):

Correct. Now do the same thing on the right hand side. You are trying to determine\[\log 7^{x-1} = ?\]Applying the same rule, what do you get?

OpenStudy (anonymous):

x-1=log 7

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Not exactly. In this one, a = 7 and b = x-1. Try again

OpenStudy (anonymous):

7 log x-1

OpenStudy (anonymous):

Backwards. Try again

OpenStudy (anonymous):

x-1 log 7??

OpenStudy (anonymous):

i got -7.74293 as an answer

OpenStudy (anonymous):

That's it. So now you done\[\log 3^{2x} = \log 7^{x-1}\]\[2x \log 3 = \left( x-1 \right) \log 7\]OK so far?

OpenStudy (anonymous):

yes! I am good

OpenStudy (anonymous):

Great. Now expand the right hand side.

OpenStudy (anonymous):

0.47712125472(2x)

OpenStudy (anonymous):

That's the left hand side. You can multiply the 0.477... by the 2.

OpenStudy (anonymous):

0.95424250943 sorry

OpenStudy (anonymous):

Excellent. The left hand side is 0.95424250944 x. Now, on to the left hand side. First thing to do is to expand it.

OpenStudy (anonymous):

** right hand side. Sorry

OpenStudy (anonymous):

0.84509804001x-0.845098804001

OpenStudy (anonymous):

right??

OpenStudy (anonymous):

Exactly well done! So now you have\[0.95424250944 x = 0.84509804001 x - 0.8450904001\]Can you gather up the x's on one side and solve?

OpenStudy (anonymous):

-0.36797678529=-0.8450904001

OpenStudy (anonymous):

so would the answer be 7.74293?

OpenStudy (anonymous):

Problem with the left hand side. Remember, you are subtracting 0.84509804001 x from both sides. Try the left hand side again.

OpenStudy (anonymous):

0.10914370543x

OpenStudy (anonymous):

That's better. So you have\[0.10914370543 x = 0.84509804001\]To solve for x, divide both sides by 0.10914370543. What do you get?

OpenStudy (anonymous):

7.74298468868 !!!!

OpenStudy (anonymous):

Yayyyy!! Well done!

OpenStudy (anonymous):

Yayyyy! If I post another question in the open section, will u answer?

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

Yup. You're welcome

OpenStudy (anonymous):

To avoid dealing with all those decimal places, most folks will leave the logs until the end, for example\[2x \log 3 = (x-1)\log 7\]\[(2 \log 3) x = (\log 7) x - \log 7\]\[(2 \log 3 - \log 7) x = -\log 7\]\[x = \frac{ -\log 7 }{ 2 \log 3 - \log 7 }\]Then plug it into your calculator